Plot a scatterplot with at least 5 data point, given that the correlation coefficient.

1
Expert's answer
2021-11-10T11:32:01-0500

Scatterplot:



Formula to calculate correlation coefficient is as follows:

"r=\frac{n(\Sigma XY)-(\Sigma X)(\Sigma Y)}{\sqrt{[n\Sigma X^{2}-(X)^{2}][n\Sigma Y^{2}-(Y)^{2}]}}"

The calculation of correlation coefficient is shown in table below:


"r=\frac{8(832)-(43)(132)}{\sqrt{[(8*281)-(43*43)][8*2994)-(132*132)]}}=0.9758"


Correlation coefficient is close to 1 that means a strong and positive association exists between two variables.